When methane (CH4) is burned in oxygen, carbon dioxide and water vapor are formed.

When 8.00 L of methane at 125°C and 1.00 atm is burned in an excess of oxygen, the

products are collected into a separate 10.0 L flask.


(a) Write a balanced equation for the reaction.

(b) What is the total pressure of the products in the flask at 125°C?

(c) What is the partial pressure of each of the products in the flask?


1
Expert's answer
2021-01-25T04:30:20-0500

(a).CH4(g)+2O2(g)"\to" CO2(g) +2H2O(g)

P1V1=P2V2

8*1=10P2

(b).P2=0.8atm

n=PV/RT

=1*8/(398*0.0820574)

=0.24496moles

Moles of CO2=0.24496moles

Moles of H2O=2*0.24496moles=0.48991moles

(c).PCO2=0.24496*0.8/(0.48991+0.24496)

=0.2667

PH2O=0.48991*0.8/0.734873

=0.5333